# AGRAG: Advanced Graph-based Retrieval-Augmented Generation for LLMs

### v1

@@ 6   Let $c^{H}_{(u,v)}\leftarrow\sum_{e\in P_{(u,v)}}c_{e}$, where $c_{e}\in\mathcal{C}_{\mathcal{E}}$;
8Compute minimum cost Steiner tree $\mathcal{G}^{\prime}$ of $H$;
@@ 9 Let $\mathcal{V}_{\text{MCMI}}\leftarrow\emptyset$, $\mathcal{E}_{\text{MCMI}}\leftarrow\emptyset$;
13Calculate $r_{\text{MCMI}}$ with Equation 16;
14 while the node $v$ in MCMI’s 1-hop neighborhood with highest $s_{v}/c_{e_{(\text{MCMI},v)}}$ satisfies $s_{v}/c_{e_{(\text{MCMI},v)}}>r_{\text{MCMI}}$ do
15   Recalculate $r_{\text{MCMI}}$ with Equation 16;
@@ III-D Step 3: Hybrid Retrieval
$\mathcal{T}_{\text{HR}}=\left\{t\in\mathcal{T}_{c}\,\middle|\,f\in\arg\top_{k_{r}}\left(\text{HS}({q},{t})\right)\right\},$ (18)
@@
where $k_{r}$ denotes the hyperparameter that controls the number of hybrid retrieved text chunks, t, q denotes the vector representation of text chunk $t$ and query $q$, respectively; $\text{HS}(q,t)$, $\text{BM25}(q,t)$ and $\text{MS}(\textbf{q},\textbf{t})$ denotes the hybrid retrieval score, the BM25 [38] sparse retrieval score, and the dense representation vectors’ cosine similarity score between query $q$ and text $t$ in Equation 7.
After obtaining the textual MCMI subgraph string $g$ and the chunks $\mathcal{T}_{\text{HR}}$ from hybrid retrieval, we input them together with the query $q$ and the the query answer generation LLM instruction $p_{\text{gen}}$, to ask the LLM to give an answer to query $q$ based on these provided contents.
@@ III-E1 Time Complexity
For the entity extraction procedure, the time complexity would be $O(|\mathcal{V}||\mathcal{T}_{c}|)$; Adding passage nodes also requires $O(|\mathcal{V}||\mathcal{T}_{c}|)$ time. For graph weighting, calculate the PPR score requires $O(|\mathcal{T}_{c}|(|\mathcal{V}|+|\mathcal{E}|))$ time, and calculate the edge cost requires $O(|\mathcal{E}|)$ time. For AGRAG’s Step 3, the time complexity of AGRAG’s MCST generation is $O(|\mathcal{E}|+|\mathcal{V}|log|\mathcal{V}|)$ using the Mehlhorn algorithm [71]; The expansion from MCST to MCMI would cost $O(|\mathcal{V}_{\text{MCMI}}|\cdot|\mathcal{N}(\mathcal{V}_{\text{MCMI}})|)$ Hence, the time complexity of AGRAG’s MCMI generation at the worst case when $\mathcal{V}_{\text{MCMI}}=\mathcal{V}$ and the graph is fully connected would be: $O(|\mathcal{V}|^{2}+|\mathcal{E}|).$
Hence, the total time complexity of AGRAG is: $O(|\mathcal{T}_{c}|(|\mathcal{V}|+|\mathcal{E}|)+|\mathcal{V}|^{2})$. Since the number of text chunk $|\mathcal{T}_{c}|$ is less or equal to the number of nodes $|\mathcal{V}|$ in graph, in the worst case, the total time complexity of AGRAG is:
@@ III-E2 Space Complexity
The graph of AGRAG’s Step 1 and Step 2 is stored with the dense weighted adjency matrix, which needs $O(|\mathcal{E}|)$ space; Storing the text chunk would need $O(|\mathcal{T}_{c}|)$ space; Storing the entity nodes require $O(|\mathcal{V}|)$ space. Hence the total space complexity of AGRAG’s Graph Indexing step would be: $O(|\mathcal{V}|+|\mathcal{E}|+|\mathcal{T}_{c}|)$. In AGRAG’s Step 3, we only need $O(|\mathcal{E}|)$ to store the graph.
@@ IV Experiments
In this section, we present the experimental evaluation of our framework AGRAG against 6 effective baselines spanning 3 technical categories across 5 tasks. Specifically, we first outline the experimental setup in Section IV-A, including details on datasets, and evaluation metrics. Next, we report the experimental results, focusing on both effectiveness and efficiency evaluations, in Section IV-B and Section IV-C, respectively. Finally, we conduct ablation study in subsection IV-D, parameter sensitivity study in subsection IV-E, and case study in subsection IV-D.
@@ IV-A1 Datasets & Tasks
We select the Novel and Medical dataset from GraphRAG-bench benchmark [34] to evaluate AGRAG’s performance on 4 tasks that have different complexity levels, these tasks have the same input and output format with the classic QA task:
@@
  • Complex Reasoning (Complex Rea.): The level 2 task of GraphRAG-bench, to solve this task, the model needs to connect multiple facts on the graph to form multi-hop, long-ranged reasoning chains.
@@
  • Text Classification: Give a set of texts $\mathcal{T}$ and a set of label $\mathcal{L}$, the classification task aim to assign each text $t\in\mathcal{T}$ a label $l\in\mathcal{L}$ that best suited to describe its semantic meaning. For text classification task, we select the Web-Of-Science (WOS) dataset [72], and the IFS-Rel dataset [73].
We select one text passage per label from the training sets of the WOS and IFS-Rel datasets to construct the corpus for RAG. During retrieval, instead of retrieving the full text chunk, we enable HippoRAG2 and AGRAG to directly retrieve the corresponding label of each text chunk. The retrieved labels are then used to form a candidate label set for the LLM’s text prediction. The LLM is subsequently tasked with selecting the most appropriate label from this candidate set.
The detailed dataset and task statistics are shown in Table II.
TABLE II: Statistics of the Novel and Medical dataset from GraphRAG-bench [34] across 4 tasks, and statistics of the WOS [72] and IFS-REL [73] dataset for the text classification task.
Dataset Creative Gen. Contextual Sum. Complex Rea. Fact Retrieval
# of Questions
Novel 67 362 289 971
Medical 166 289 509 1,098
Dataset Text Classification
# of Texts # of Labels Token Per Text
WOS 9,396 133 200
IFS-REL 2,560 64 105
@@
Community
-based
@@
Graph-
wandering
HippoRAG2 0.496 0.295 0.325 0.347 0.563 0.451 0.735 0.586 0.496 0.560 0.314 0.453 0.332 0.622 0.310 0.470 0.343 0.598 0.424 0.270 0.389 0.246
@@
TABLE IV: The performance comparasion of AGRAG and baselines with GPT-4o [74] and R1-Searcher-Qwen-2.5-7B [75] backbone
Backbone Task Creative Gen. Contextual Sum. Complex Rea. Fact Retrieval
Dataset Novel Medical Novel Medical Novel Medical Novel Medical
GPT-4o HippoRAG2 0.519 0.429 0.355 0.322 0.571 0.545 0.752 0.612 0.671 0.597 0.340 0.477 0.391 0.667 0.340 0.490 0.400 0.631
AGRAG 0.536 0.478 0.414 0.374 0.632 0.609 0.794 0.667 0.703 0.634 0.351 0.515 0.397 0.673 0.343 0.503 0.402 0.649
R1-Searcher NaiveRAG 0.211 0.203 0.255 0.210 0.036 0.124 0.604 0.444 0.181 0.254 0.069 0.285 0.050 0.200 0.086 0.306 0.063 0.208
HippoRAG2 0.336 0.187 0.259 0.275 0.258 0.250 0.708 0.483 0.456 0.416 0.184 0.304 0.231 0.448 0.215 0.272 0.242 0.490
AGRAG 0.400 0.255 0.263 0.308 0.270 0.261 0.715 0.495 0.480 0.429 0.208 0.326 0.245 0.464 0.313 0.303 0.253 0.511
@@ IV-A2 Evaluation Metrics
In this paper, we apply 5 metrics to evaluate AGRAG’s performance on 5 different tasks:
@@
    ROUGE-L (ROG.) [76]: Evaluates the answer quality of by comparing the longest common subsequence (LCS) between the generated answer $G$ and the gold answer $Y$.
    $\text{ROUGE-L}=\frac{\text{LCS}(G,Y)}{|G|},$ (21)
    we apply ROUGE-L to the fact retrieval and the complex reasoning tasks, as the query answer of these two tasks are usually a fixed entity or relation.
@@
    Coverage (COV.) [34]: Evaluates answer completeness by assessing whether all necessary reference evidences $E$ are included in the LLM’s response $R$, computed as the proportion of covered evidences to the number of required evidences $|E|$:
    $\text{Cov}=\frac{|\{e\in E\ |\ M(e,R)\}|}{|E|},$ (22)
    where $M(e,R)$ is a boolean function indicating whether evidence $e$ is covered by generated response $R$. We apply the coverage metric to contextual summarization and creative generation tasks, as the their answers are usually passages that conclude multiple information points.
@@
    Faithfulness (FS.) [34]: Measures the percentage of answer claims $A$ directly supported by retrieved context $C$, calculated as the ratio of claims aligned with evidence to the number of total claims $A$:
    $\text{FS}=\frac{|\{a\in A\ |\ S(a,C)\}|}{|A|},$ (23)
    where $S(a,C)$ is a boolean function indicating whether a claim is supported by contexts in $C$. We apply the faithfulness metric to the creative generation task to evaluate the quality of the hypothesis in the LLM’s response.
TABLE V: Model efficiency on 4 datasets for 5 tasks, all experiments were carried out with Qwen2.5-14B-Instruct LLM backbone. Where AGRAG achieve at most 166.3% accleration, and 368.8% less token cost compared with tested baselines.
Category Models Time Cost (mins) Token Cost (# per query)
@@
Traditional
RAG
NaiveRAG 384.5 (+85%) 252.3  (+63%) 294.2 (+98%) 227.1 (+64%) 11,145 (+46%) 7,852 (+41%) 7,321 (+80%) 4,537 (+129%)
L.LLMLingua 410.5 (+98%) 342.8 (+122%) 362.2 (+144%) 318.7 (+13%) 10,216 (+34%) 6,575 (+18%) 5,892 (+45%) 2,771 (+40%)
Community
-based
GraphRAG 552.6 (+166%) 352.1 (+128%) 325.0 (+119%) 281.7 (+104%) 14,795 (+94%) 11,511 (+107%) 12,160 (+199%) 9,296 (+369%)
LightRAG 386.5 (+86%) 279.5 (+81%) 252.7 (+70%) 229.4 (+66%) 14,377 (+89%) 9,961 (+79%) 11,737 (+188%) 7,866 (+297%)
Graph-
wandering
HippoRAG2 229.2 (+11%) 176.1 (+14%) 190.4 (+28%) 154.7 (+12%) 8,978 (+18%) 6,859 (+23%) 5,159 (+27%) 2,542 (+28%)
F.GraphRAG 319.8 (+54%) 221.44 (+43%) 215.4 (+45%) 191.1 (+38%) 8,742 (+15%) 6,076 (+9%) 4,992 (+23%) 2,237 (+13%)
AGRAG 207.5 154.4 148.6 138.4 7,627 5,558 4,071 1,983
@@ IV-A3 Baselines
We compare AGRAG’s performance with 6 baselines from 3 technical categories, they are traditional RAG model NaiveRAG [15] and LongLLMLingua [69]; Community-based model GraphRAG [25], and LightRAG [26]; And walking-based model Fast-GraphRAG [32] and HippoRAG2 [31]. We use the Qwen-2.5-14B-Instruct [77] as the LLM backbone unless further illustrated.
Traditional RAG Models These models directly retrieve text chunks based on the semantic similarity between chunks and the query.
@@
  • HippoRAG2 [31]: It employs LLM for graph construction, and also add the text chunks as nodes in the graph. As a result, it directly retrieves Top-k text chunks based on their respective nodes’ PPR score.

### v2

@@ 6   Let $c^{H}_{(u,v)}\leftarrow\sum_{e\in P_{(u,v)}}c_{e}$, where $c_{e}\in\mathcal{C}_{\mathcal{E}}$;
8Compute an approximated MCST $\mathcal{G}^{\prime}$ of $H$;
@@ 9 Let $\mathcal{V}_{\text{MCMI}}\leftarrow\emptyset$, $\mathcal{E}_{\text{MCMI}}\leftarrow\emptyset$;
13Calculate $r(\mathcal{G}_{\text{MCMI}})$ with Equation (15);
14 while the node $v$ in MCMI’s 1-hop neighborhood with lowest $c_{e_{(\text{MCMI},v)}}/s_{v}$ satisfies $c_{e_{(\text{MCMI},v)}}/s_{v}<r(\mathcal{G}_{\text{MCMI}})$ do
15   Recalculate $r(\mathcal{G}_{\text{MCMI}})$ with Equation (15);
@@ III-D Step 3: Hybrid Retrieval
$\mathcal{T}_{\text{HR}}=\left\{t\in\mathcal{T}_{c}\,\middle|\,t\in\arg\top_{k_{r}}\left(\text{HS}({q},{t})\right)\right\},$ (20)
@@
where $k_{r}$ denotes the hyperparameter that controls the number of hybrid retrieved text chunks, t, q denotes the vector representation of text chunk $t$ and query $q$, respectively; $\text{HS}(q,t)$, $\text{BM25}(q,t)$ and $\text{MS}(\textbf{q},\textbf{t})$ denotes the hybrid retrieval score, the BM25 [59] sparse retrieval score, and the dense representation vectors’ cosine similarity score between query $q$ and text $t$ in Equation (7).
After obtaining the textual MCMI subgraph string $g$ and the chunks $\mathcal{T}_{\text{HR}}$ from hybrid retrieval, we input them together with the query $q$ and the query answer generation LLM instruction $p_{\text{gen}}$, to ask the LLM to give an answer to query $q$ based on these provided contents.
@@ III-E1 Time Complexity
For the entity extraction procedure, the time complexity is $O(|\mathcal{V}||\mathcal{T}_{c}|)$; Adding passage nodes also requires $O(|\mathcal{V}||\mathcal{T}_{c}|)$ time. For graph weighting, calculating the PPR score requires $O(|\mathcal{T}_{c}|(|\mathcal{V}|+|\mathcal{E}|))$ time, and calculate the edge cost requires $O(|\mathcal{E}|)$ time. For AGRAG’s Step 3, the time complexity of AGRAG’s MCST generation is $O(|\mathcal{E}|+|\mathcal{V}|log|\mathcal{V}|)$ using the Mehlhorn algorithm [50]; The expansion from MCST to MCMI would cost $O(|\mathcal{V}_{\text{MCMI}}|\cdot|\mathcal{N}(\mathcal{V}_{\text{MCMI}})|)$ Hence, the time complexity of AGRAG’s MCMI generation at the worst case when $\mathcal{V}_{\text{MCMI}}=\mathcal{V}$ and the graph is fully connected is: $O(|\mathcal{V}|^{2}+|\mathcal{E}|).$
Hence, the end to end time complexity of AGRAG is: $O(|\mathcal{T}_{c}|(|\mathcal{V}|+|\mathcal{E}|)+|\mathcal{V}|^{2})$. Since the number of text chunk $|\mathcal{T}_{c}|$ is less or equal to the number of nodes $|\mathcal{V}|$ in graph, in the worst case, the end to end time complexity of AGRAG is:
@@ III-E2 Space Complexity
The graph of AGRAG’s Step 1 and Step 2 is stored with the dense weighted adjacency matrix, which needs $O(|\mathcal{E}|)$ space; Storing the text chunk would need $O(|\mathcal{T}_{c}|)$ space; Storing the entity nodes require $O(|\mathcal{V}|)$ space. Hence the total space complexity of AGRAG’s Graph Indexing step is: $O(|\mathcal{V}|+|\mathcal{E}|+|\mathcal{T}_{c}|)$. In AGRAG’s Step 3, we only need $O(|\mathcal{E}|)$ to store the graph.
@@ IV Experiments
In this section, we first outline the experimental setup in Section IV-A, including details on datasets, and evaluation metrics. Next, we report the experimental results, focusing on both effectiveness and efficiency evaluations, in Section IV-B and Section IV-C, respectively. Finally, we conduct ablation study in Section IV-D, and parameter sensitivity study in Section IV-E.
@@ IV-A1 Datasets & Tasks
We select the Novel and Medical datasets from GraphRAG-bench [74] to evaluate AGRAG’s performance on 4 tasks that have different complexity levels:
@@
  • Complex Reasoning (Complex Rea.): The level 2 task of GraphRAG-bench, to solve this task, the model needs to connect multiple facts on the graph to form multi-hop, long-range reasoning chains.
@@
  • Text Classification: Given a set of texts $\mathcal{T}$ and a set of label $\mathcal{L}$, the classification task aim to assign each text $t\in\mathcal{T}$ a label $l\in\mathcal{L}$ that best suited to describe its semantic meaning. For text classification task, we select the Web-Of-Science (WOS) dataset [38], and the IFS-Rel dataset [73].
We select one text passage per label from the training sets of the WOS and IFS-Rel datasets to construct the corpus for RAG. During retrieval, instead of retrieving the full text chunk, we enable HippoRAG2 and AGRAG to directly retrieve the corresponding label of each text chunk. The retrieved labels are then used to form a candidate label set for the LLM’s text prediction. The LLM is subsequently tasked with selecting the most appropriate label from this candidate set. The detailed dataset and task statistics are shown in Table II.
Table II: Statistics of the Novel and Medical dataset from GraphRAG-bench [74] across 4 tasks, and statistics of the WOS [38] and IFS-REL [73] dataset for the text classification task. CG, CS, CR, FR and TC denotes the Creative Generation, Contextual Summarization, Complex Reasoning, Fact Retrieval, and Text Classification task, respectively. ‘-’ denotes the respective task is not included by certain dataset.
Dataset Chunks KG Entities KG Relations AVG MCST Nodes AVG MCST Edges AVG MCMI Nodes AVG MCMI Edges CG. CS. CR. FR. TC.
Questions Texts Labels
Novel 5,140 63,887 163,169 8.7 10.7 19.7 35.9 67 362 610 971 - -
Medical 1,030 12,426 52,054 8.9 9.9 23.7 59.5 166 289 509 1,098 - -
WOS 24 475 923 15.5 17.8 23.8 35.8 - - - - 9,396 133
IFS-Rel 11 185 355 11.0 12.8 17.0 26.3 - - - - 2,560 64
@@
Tree -based RAG RAPTOR 0.365 0.333 0.373 0.314 0.526 0.553 0.760 0.626 0.633 0.583 0.290 0.460 0.185 0.649 0.133 0.465 0.132 0.609 0.409 0.290 0.293 0.212
SiReRAG 0.324 0.328 0.365 0.295 0.544 0.456 0.754 0.617 0.592 0.560 0.307 0.459 0.269 0.646 0.311 0.451 0.349 0.558 0.430 0.294 0.325 0.220
Community -based RAG G-retriever 0.294 0.167 0.200 0.169 0.371 0.468 0.594 0.404 0.302 0.395 0.146 0.441 0.102 0.396 0.199 0.209 0.148 0.326 0.188 0.129 0.311 0.185
MIXRAG 0.200 0.286 0.337 0.255 0.351 0.306 0.695 0.538 0.409 0.464 0.109 0.386 0.156 0.434 0.102 0.366 0.160 0.463 0.244 0.235 0.362 0.203
@@
Walking -based RAG HippoRAG2 0.496 0.295 0.325 0.347 0.563 0.451 0.735 0.586 0.496 0.560 0.314 0.453 0.332 0.622 0.310 0.470 0.343 0.598 0.424 0.270 0.389 0.246
@@ IV-A2 Evaluation Metrics
In this paper, we apply 5 metrics with the RAGAS [68] package to evaluate AGRAG’s performance on 5 different tasks:
@@
    ROUGE-L (ROG.) [44]: Evaluates the answer quality by comparing the longest common subsequence (LCS) between the generated answer $G$ and the gold answer $Y$: $\text{ROUGE-L}=\frac{\text{LCS}(G,Y)}{|G|}$. We apply ROUGE-L to the fact retrieval and the complex reasoning tasks, as the answer of these two tasks are usually a fixed entity or relation.
@@
    Coverage (COV.) [74]: Evaluates answer completeness by assessing whether all necessary reference evidences $E$ are included in the LLM’s response $R$, computed as the proportion of covered evidences to the number of required evidences $|E|,\text{Cov}=\frac{|\{e\in E\ |\ M(e,R)\}|}{|E|}$, where $M(e,R)$ is a boolean function indicating whether evidence $e$ is covered by generated response $R$. We apply the coverage metric to contextual summarization and creative generation tasks, as their answers are usually passages that conclude multiple information points.
@@
    Faithfulness (FS.) [74]: Measures the percentage of answer claims $A$ directly supported by retrieved context $C$, calculated as the ratio of claims aligned with evidence to the number of total claims $A$: $\text{FS}=\frac{|\{a\in A\ |\ S(a,C)\}|}{|A|},$ where $S(a,C)$ is a boolean function indicating whether a claim is supported by contexts in $C$. We apply the faithfulness metric to the creative generation task to evaluate the quality of the hypothesis in the LLM’s response.
Table IV: Model efficiency on 4 datasets for 5 tasks, all experiments were carried out with Qwen2.5-14B-Instruct LLM backbone.
Category Models Time Cost (Indexing / Querying mins) Token Cost (Input / Output per query)
@@
Traditional RAG NaiveRAG 13 / 371 7 / 245 2 / 292 2 / 225 9,437 / 1,708 7,473 / 379 7,292 / 29 4,521 / 16
L.LLMLingua 51 / 360 37 / 305 4 / 358 3 / 315 8,466 / 1,750 5,686 / 390 5,861 / 31 2,757 / 14
Tree -based RAG RAPTOR 464 / 250 117 / 319 30 / 162 24 / 121 7,243 / 777 4,817 / 695 6,007 / 41 3,807 / 33
SiReRAG 85 / 233 40 / 207 26 / 139 18 / 117 7,759 / 800 4,918 / 701 6,012 / 47 3,821 / 36
Community -based RAG G-Retriever 19 / 119 12 / 145 8 / 158 3 / 164 6,344 / 116 6,290 / 157 6,762 / 12 4,002 / 21
MIXRAG 17 / 388 11 / 453 5 / 257 2 / 193 6,083 / 237 4,238 / 230 4,193 / 31 3,597 / 29
GraphRAG 347 / 205 194 / 158 85 / 240 64 / 217 12,686 / 2,109 8,385 / 1,017 11,809 / 351 9,009 / 287
LightRAG 166 / 221 116 / 164 42 / 211 31 / 197 13,145 / 1,233 9,535 / 426 11,510 / 227 7,667 / 199
Walking -based RAG HippoRAG2 95 / 134 46 / 130 19 / 140 15 / 127 8,138 / 1,073 6,264 / 595 5,066 / 93 2,440 / 102
F.GraphRAG 106 / 198 48 / 173 29 / 186 27 / 164 7,490 / 1,252 6,475 / 384 4,957 / 35 2,217 / 20
AGRAG 65 / 142 16 / 138 16 / 132 14 / 124 6,487 / 1,140 5,145 / 413 3,970 / 101 1,890 / 93
@@ IV-A3 Baselines
We compare AGRAG’s performance with 6 baselines from 3 technical categories, they are traditional RAG model NaiveRAG [22] and LongLLMLingua [34]; Community-based model GraphRAG [17], and LightRAG [24]; And walking-based model Fast-GraphRAG [13] and HippoRAG2 [gutiérrez2025ragmemory]. We use the Qwen-2.5-14B-Instruct [57] as the LLM backbone unless further illustrated.
Traditional RAG Models These models directly retrieve text chunks based on their semantic similarity to the queries.
@@
Tree-based Models These models organize text chunks into tree structures to capture long-range dependencies and thematic information.
  • RAPTOR [60] RAPTOR recursively clusters and summarizes text chunks to construct a tree from the bottom up, enabling the retrieval of information at varying levels of abstraction to answer high-level thematic queries.
  • SiReRAG [82] Based on RAPTOR, SiReRAG builds similarity and relatedness trees to index respective information, thereby supporting complex multihop reasoning.
@@
  • G-retriever [28] It formulates subgraph retrieval as a Prize-Collecting Steiner Tree (PCST) optimization problem, identifying connected subgraphs that maximize relevance while minimizing costs.
  • MIXRAG [45] It introduces a Mixture-of-Experts (MoE) framework to handle diverse query intents, dynamically routing queries to specialized experts and employing a query-aware graph encoder to filter noise.
@@
  • HippoRAG2 [gutiérrez2025ragmemory]: It employs LLM for graph construction, and also add the text chunks as nodes in the graph. As a result, it directly retrieves Top-k text chunks based on their respective nodes’ PPR score.